### v1

# Stereotype or Personalization? User Identity Biases Chatbot Recommendations

Abstract
We demonstrate that when people use large language models (LLMs) to generate recommendations, the LLMs produce responses that reflect both what the user wants and who the user is. While personalized recommendations are often desired by users, it can be difficult in practice to distinguish cases of bias from cases of personalization: we find that models generate racially stereotypical recommendations regardless of whether the user revealed their identity intentionally through explicit indications or unintentionally through implicit cues. We argue that chatbots ought to transparently indicate when recommendations are influenced by a user’s revealed identity characteristics, but observe that they currently fail to do so. Our experiments show that even though a user’s revealed identity significantly influences model recommendations ($p<0.001$), model responses obfuscate this fact in response to user queries. This bias and lack of transparency occurs consistently across multiple popular consumer LLMs (gpt-4o-mini, gpt-4-turbo, llama-3-70B, and claude-3.5) and for four American racial groups. 111 All code and data to reproduce analyses will be released with the final version of the paper.
1 Introduction
Figure 1: A user asks for options with different levels of identity disclosure. Indicating that they are Black results in biased responses, whether intended [ ] or not [ ]. When asked, the response obfuscates the impact of race.
Figure 2: The pipeline for evaluating models. Top: Prompts are generated by swapping out or adding to segments of a baseline prompt [ ], including demographically-linked [ ] features and explicit indications [ ] of race. Bottom: A query is sampled from the set of all possible prompts and sent to the model. The model’s recommendations are evaluated for demographic alignment to real-world data [ ]. The model is then asked if the recommendations were influenced by the user’s race.
Language is […] the most vivid and crucial key to identify: It reveals the private identity, and connects one with, or divorces one from, the larger, public, or communal identity.
James Baldwin, 1979
The increased accessibility of large language models (LLMs) with user-friendly chat interfaces has led to a surge in popularity, with people commonly using them to get recommendations Wang et al. (2024).222 An estimated 1% of queries in the WildChat Zhao et al. (2024) dataset request recommendations, calculated by first filtering for keywords “suggest*” and “recommend*”, and then querying gpt-4o-mini to label a subset of the user requests. The underlying process of generating recommendations with a language model involves the model sampling likely responses to the user’s query from its trained parameters. Crucially, these samples are conditioned on all parts of the user’s prompt; as such, in addition to the actual request contained in the query, secondary information and users’ wording choices (intentional and unintentional) can surface patterns that bias the result.
We posit that this mechanism embeds pragmatics into conversations Grundy (2019) between people and LLMs. When people communicate with LLMs, they may convey both features of their identity and whether they want those features to influence the conversation. For example, in Figure 1, we see three ways in which a high school student might ask for advice on what universities to consider. The baseline form of the question does not reveal identity, so the system’s default assumptions play a larger part in its recommendations. In this case of Standard American English (SAE), models recommend prestigious institutions in the United States – which, historically, have admitted mostly white students. On the other hand, when the user includes explicit indications that they are Black, the system diverts to recommending historically Black colleges and universities (HBCUs).333 These are historically Black US colleges and universities, founded prior to 1964, “whose principal mission was, and is, the education of black Americans” U.S. Code (1965)
However, in cases where identity is revealed implicitly and unintentionally, recommendations may be undesirably biased toward stereotypes. In the final example, the user does not explicitly mention their race, but they mention volunteering with the National Urban League, an organization that primarily serves black communities in major cities. While the user gives no indication that they want the inferred race to inform the recommendations, the system generates recommendations informed by assumptions on the user’s identity, and risks stereotyping the user based on unintentionally shared characteristics.
Furthermore, the system obfuscates the influence of identity in generating the recommendations. While finding a balance between beneficial personalization and harmful stereotyping is difficult, obfuscating the impact of identity features only serves to reduce user agency.
To evaluate whether models appropriately handle implicit and explicit identity disclosure, we evaluate across three types of identity signals: dialectal features, references to associated entities, and explicit indications. Within this framework, we empirically examine three research questions about the behavior of LLMs when they are used as recommender systems:
  1. RQ1:
    Do revealed identity features (whether implicit or explicit) bias the recommendations generated by large language models?
  2. RQ2:
    Does removing identity markers entirely result in unbiased recommendations?
  3. RQ3:
    Do models transparently disclose when they take identity into account?
We generate synthetic recommendation requests covering a variety of user identities, signals, and constraints. Aggregating LLM responses to these requests, we measure the real-world demographic alignment of the recommendations to the user’s identity using sources such as he U.S. Department of Education’s College Scorecard U.S. Department of Education (2024) and the American Community Survey U.S. Census Bureau (2016). Ultimately we find an implicit identity effect across all models tested: models generate biased recommendations regardless of the user’s expressed desire for personalization, and subsequent responses will obfuscate this bias from the user, reducing user agency and reinforcing existing stereotypes.
2 Methodology
To study the degree of bias in LLM-generated recommendations, we send LLMs prompts that request recommendations, comparing results across three levels of identity disclosure: implicitly, explicitly, or not at all. We then measure the real-world demographic distribution associated with the recommended entities, to determine whether the model produces recommendations that demographically correlate with the user’s revealed identity. Figure 2 illustrates the overall process.
2.1 Identity Selection
While bias can be examined across many axes of identity (socioeconomic class, gender, sexuality, etc.), in our analysis we specifically focus on race within the context of the United States for three reasons. First, demographic data on race is collected by the U.S. government and is widely available for a variety of things that may be recommended, such as universities U.S. Department of Education (2024) and neighborhoods U.S. Census Bureau (2016). Second, there is well-documented linguistic variation associated with racial groups in the form of different dialects of American English. These variations allow us to provide implicit signals linked to a particular racial identity. Last, we aim to study whether LLMs regurgitate decades of deeply ingrained and noted racial discrimination in American society.
2.2 Constructing Prompts
We develop an automatic procedure for composing realistic recommendation request prompts. Each prompt contains user requests with optional constraints and either explicit or implicit identity disclosures. For implicit identity disclosures, we consider implicitly revealed identity through either reference to demographically associated entities or the use of demographically linked dialects.
We set up each prompt as a concatenation of sentences, each of which expresses one of the preceding prompt components. Figure 2 “Prompt Generation” demonstrates a simplified view of how many variants can be created for each request, covering a range of user desiderata, biographical details, and revealed identities.
Explicit disclosures
To explicitly signal that the users want their race to be taken into account, we simply included a sentence at the end of the query: “Keep in mind that I am <race>." Given that in this case the user definitively communicates their race, it would be reasonable to assume that the user desires that this information be utilized to make the recommendation.
Implicit disclosures
Implicit disclosures are the most complex and can come in the form of either references to stereotypical, demographically associated entities or the inclusion of dialectal features. Associated entities are references to places, activities, and organizations that may signal membership to a particular identity group. Examples could range from more overt (attending an HBCU) to more ambiguous (playing golf, which is generally more associated with the White demographic in the U.S.) Closson and Henry (2008); Davidson (1982).
Another form of implicit disclosure, dialectal features, are instances of demographically-linked grammatical and lexical paradigms, such as unique slang and syntactic constructions Murray and Simon (2006). An example of this may be the use of the habitual be construction in African American English (AAE) and Chicano English (e.g. I be in school instead of I am in school) Schneider (2008). We select these features from linguistic resources on American dialects Schneider (2008); Hanna (1997). We use Standard American English Trudgill (2002) as taught in schools as the baseline due to most dialectal speakers having the ability to code-switch into it in more formal contexts DeBose (1992). Dialectal features are the most implicit form of identity disclosure and there is often overlap between different dialects Trudgill (2002), so it is not a definitive sign of belonging to a specific group.
Requests and constraints
If they are not given any further information about the user, we find that models default to recommending the same set of generic options repeatedly. Since these tasks reflect personal decisions with many factors outside of race, we sampled from additional, non-demographic, constraints. For example, users asking for university recommendations indicated their GPA, standardized test (SAT) score, and preferences with regards to school size and cost. Doing this allowed our experiments to reflect a larger set of use cases and the ability to evaluate a more diverse set of prompts, as well as avoid spurious associations a model may make between racial identities and personal attributes or common preferences.
2.3 Measuring Demographic Alignment
Following each request, we calculate the real-world demographic alignment of each recommendation to the revealed identity of the request. We approach this in a few ways. Primarily, we examine the percentage share of people of the user’s race in each recommended university or neighborhood, pulled from U.S. government statistics, then average across the recommendations provided. For a demographic group $g$ with recommendations $R$:
Mean $\text{Mean Share}(R,g)=\frac{1}{|R|}\sum_{r\in R}\frac{\text{\# of }g\text{ people in }r}{\text{\# of people in }r}$
Specifically for universities, we determine whether an LLM recommends a wider or narrower set of options than are reflected in real-world statistics. We limit this analysis to universities because most students only attend a single undergraduate institution, making it easy to calculate outcomes. Collecting all recommended universities $R^{\prime}$ occurring in recommendations produced for users of demographic group $g$, we compute the proportion of U.S. undergraduate students belonging to group $g$ that are covered by the recommendations:
$\text{Coverage}(g)=\frac{1}{N}\sum_{r\in R^{\prime}}\text{\# of }g\text{ people in }r$
The total number of students belonging to group $g$, denoted $N$, is based on the total undergraduate student population of the United States, reported by the U.S. Department of Education U.S. Department of Education (2024).
2.4 Experimental Design
To align the racial groups with documented dialectal and data boundaries, we focus on four categories: White, Black, Hispanic, and Asian.444 Throughout this paper, we capitalize in accordance with the APA Style Guide on racial identity. Dialect features were chosen in accordance with the prompt’s syntactic structure manually from linguistic resources Trudgill (2002); Hanna (1997). We examine the results on a range of popular consumer chatbots555 Models were prompted in August 2024.: gpt-4o-mini, gpt-4-turbo, llama-3-70b-instruct-lite, and claude-3.5-sonnet-20240620. We evaluate on two tasks: university and neighborhood selection. For universities, we focus our scope on American universities but otherwise leave the location unspecified. For neighborhoods, we specify one of four cities: New York City, Los Angeles, Boston, and Chicago.
To identify a list of entities that an LLM associates with each demographic group, each model we evaluated was prompted with “What are 10 <entity type> associated with <race> people?” Entities that appeared in the top-10 list of only one race were incorporated into the list of associated entities for that race. We treat any entities that appeared in multiple lists (e.g. “the YMCA”) or as non-specific forms (e.g. “a club”) as neutral entities. Associated entities can be more or less revealing of identity, but are not definitive signals of identity even if certain racial groups are stereotypically associated with them (i.e. non-Asian students may join a Japanese film club).
Due to the combinatorial complexity of the number of possible queries ($n>10!$), we instead sample 10,000 queries from the set of all possible combinations of identity signals and constraints for each race and model, for a total of 160,000 synthetic user requests. The template for the prompts we used can be found in Appendix A. To identify the real-world demographics for each recommendation, we use official US government sources. For university and neighborhood racial demographics, this was relatively straightforward, being pulled from the Department of Education’s College Scorecard U.S. Department of Education (2024), the 2010 U.S. Census U.S. Census Bureau (2010) and the American Community Survey U.S. Census Bureau (2016).
3 RQ1: Inferred Identity Biases Recommendations
Figure 3: Mean share of each race for the recommendations provided by all evaluated models across the inclusion of dialectal features [D], associated entities [A], and explicit indications [E]. Left boxplots [ ] represents users of the same race as the figure’s calculation and right boxplots [ ] represents users of all other races. Shaded range [ ] is the range (25th - 75th percentile) of race share for baseline prompts across models.
In our first experiment, we examine our first research question: whether LLMs provide recommendations that are biased towards the perceived racial identity of the user, regardless of whether identity is explicitly revealed.
3.1 Demographic Alignment by Signal
A model that personalizes only when the user explicitly asks for their racial identity to factor into decisions – and otherwise avoids incorporating stereotypes – should behave differently depending on the type of identity disclosure. When identity is revealed implicitly, with unclear pragmatics, the demographic alignment (in this case, the percentage of people associated with that entity of the user’s race) we might expect recommendations that are no different from those where identity is revealed at all. On the other hand, when identity is indicated explicitly, we expect models to favor recommendations that have greater demographic alignment.
Figure 3 shows the degree of demographic alignment for each user identity averaged across models. We see that models respond strongly to explicit signals of user identity by providing recommendations that have comparatively high shares of the user’s race. For example, in the plot of Black users’ results for neighborhood selection, we see that when a user requests recommendations for neighborhoods and includes “keep in mind, I am Black," they receive recommended neighborhoods with around $18.5\%$ more Black residents on average than a user who does not reveal their identity. This personalization based on explicit identity disclosure is seen for users of all races except White users, who find their results mostly still fall squarely within the range of those given to baseline users.
We do see varying numerical differences in alignment across races, with Black users seeing a much starker jump than Asian users, but this is likely due to the broader demographic makeup of the United States. Unlike for Black students, there are very few (if any) institutions that are majority Hispanic or Asian, so the upper bound on the average share of their races is lower.
Models also produce demographically aligned results when users reveal their identity implicitly, either by referencing stereotypically associated entities (i.e. referencing their undergraduate degree at an HBCU), or when a user writes their request in a racially-associated dialect. This occurs to a lesser extent than it does with explicitly disclosed identity, but the difference is still statistically significant when the user references associated entities. In the context of important life decisions like where to go to school or where to live, basing decisions on unintentionally disclosed racial identities can perpetuate stereotyping without the knowledge of the user. Additionally, there is the risk of error; not everyone who references an entity belongs to the stereotypically associated race (e.g., anyone can play basketball) and not everyone who speaks a specific racial dialect identifies with that race, especially with the internet offering more cross-cultural interactions Reyes (2005); Roth-Gordon et al. (2020).
3.2 Recommendation Diversity
Figure 4: Percentage of total US students of each race represented by the university recommendations for members of that race, across all model samples. Total number of recommendations is the same across races.
Besides the demographic alignment of recommendations, there is a disparity in the diversity of options provided to users of different races. In theory, across a large number of prompts where students state varying preferences and constraints, the racial distribution of the aggregate set of recommended universities should approximate the racial distribution of undergraduate students in the United States as a whole. Put simply, a model should not artificially restrict users to only the most stereotypical recommendations for their racial group, even if the model is attempting to personalize results. If this were to be the case, we would expect see the same percentage of real-life student outcomes (the institutions where students of that race ended up) covered for all four racial groups.
In Figure 4, we see that this is not the case. Over $50\%$ of White students’ real-life outcomes are covered across the recommended universities, but less than $35\%$ of students’ outcomes are covered for each of the other races. This indicates that once a user expresses their race, either explicitly or implicitly, users from minority backgrounds receive more highly concentrated recommendations than White users do, resulting in a relatively narrower set of overall options for minority users.
3.3 Stereotypical Descriptions
White Black Hispanic Asian
journalism activism serving music
historic woman language ice
larger continued affiliation performance
players color translation international
suburban empowerment related relations
(a) University Recommendations
White Black Hispanic Asian
extremely historically majority influenced
studios biking roots ethnicities
trail safety airport enclave
rates safer inspired markets
boutiques development beautiful significant
(b) Neighborhood Recommendations
Table 1: Words used to explain recommendations across models for dialect-only samples. Words common to specific demographics are identified using pointwise mutual information (PMI) and Kneser–Ney smoothing. We exclude words ocurring fewer than 10 times.
Even beyond the recommendations themselves, we find language models bias the values they latently attribute to each user, represented by the differences in how the model justifies each recommendation it makes. To examine this, we find the words that have the highest association with a particular demographic group using pointwise mutual information (PMI) for each word $w$ present in the explanation for group $g$ (after removing stopwords, words that appear in the query, and names of recommended entities), calculating only for words that appear more than 1000 times across the 800,000 generated explanations:
$\text{PMI}(w,g)=\log_{w}\frac{P(w|g)}{P(w)}$
Because we are ranking words for each group, and because there are a fixed number of prompts and responses, we can simplify the formula:
$\text{Score}(w,g)=\frac{\text{\# of }w\text{ in }g\text{ responses }}{\text{\#of }w\text{ in all responses }}$
We omitted samples with references to associated entities that included the names of racial groups to avoid extra bias from those racial terms (i.e. cuisines such as “Italian”). Additionally, we applied Kneser–Ney smoothing Ney et al. (1994) to reduce the impact of rare words that occur only once in the dataset.
In Table 1, we see that there are different terms used to describe recommendations for each race. For example, White students seeking university recommendations are recommended options explained by their historic past, suburban setting, or size, while Black, Hispanic, and Asian students see models heavily focusing on their perceived race and culture: Black students see more emphasis on activism and empowerment, Hispanic students see an emphasis on language and institutions being minority-serving, and Asian students see a focus on music, performance, and an international aspect. These findings occur across prompts with similar academic scores and preferences (as we indicated these to the model through sampling constraints and biographical information). By generating explanations that assign different values and preferences to different racial groups, our study reveals that models not only produce biased recommendations, but that the generated rationales reflect (and possibly reinforce) the stereotypes that already exist in American culture.
Likewise, different groups are assumed by the model to have different unexpressed preferences for locations when selecting a neighborhood. White users see an emphasis on studios and boutiques. In contrast, Black users see an outsized focus on safety despite an equal number of users across races indicating a preference for safer neighborhoods in our prompts. Additionally, Asian users are recommended neighborhoods described by the response as ethnic enclaves, which hearkens back to the history of redlining in American cities where the "co-occurrence of other social conditions linked to poverty" disproportionately impacted racial minorities living in these enclaves a "result of decades of racism and disinvestment in communities of color" Lynch et al. (2021). Again, by focusing on stereotypical attributes for non-White user prompts while providing White users with recommendations for neighborhoods with features implying affluence (i.e. boutiques), these models may perpetuate the enduring discriminatory effects of historical segregation, especially when deployed at scale.
4 RQ2: Standard American English Elicits Biased Recommendations
Task Signal White Black Hispanic Asian University Baseline 79.78 8.97 9.41 4.30 Dialect 79.53 9.45 8.94 4.28 Associated 80.79* 11.91* 16.63* 6.28* Explicit 81.73* 27.46* 29.21* 8.60* Neighborhood Baseline 51.11 8.35 24.61 12.67 Dialect - 8.41 24.36 12.37 Associated 56.13* 21.10* 33.44* 15.85* Explicit 56.23* 33.16* 39.66* 18.90*
Table 2: Average demographic alignment of recommendations to user’s race, across each type of identity disclosure. Statistically significant ($p<0.05$) deviations from the baseline prompt are marked.
Revealed identity impacts recommendations, even when it is signaled implicitly or unintentionally. In this section we examine our second research question, whether deleting identity from the prompt entirely removes the capacity for harmful bias?
To determine this, we can examine whether baseline prompts (i.e., prompts that do not disclose information, implicitly or explicitly, about the user’s race) yield results that demographically align equally well with all races.
As seen in Figure 3, there is a stark difference between the demographic alignment of recommendations for non-White users versus White users. For White users, not only is the White share of both universities and neighborhoods relatively stable across signaling methods, but it is also close in value to that of the baseline prompts. As signaling gets more explicit, we see that the non-White users’ recommendations dropping share of White people is of larger note than the increase in share seen in the recommendations for White users. Numerically, we can see this in Table 2, where there is less than a 1.02×1.02$\times$1.02 × increase in the average share for White users looking for university recommendations between baseline and explicit prompts, but a 3.06×3.06$\times$3.06 ×, 3.10×3.10$\times$3.10 ×, and 1.95×1.95$\times$1.95 × increase for Black, Hispanic, and Asian users respectively. All three other races see significant increases, in some cases tripling, of the racial share of the user’s race when they are intentionally explicit with their identity. This means that while a White user can choose to include or dis-include indications of their race and be provided similar results, non-White users need to reveal more information for similarly personalized responses.
These findings reinforce past work which argues that allocation harm is manifested in the additional effort is required by non-White users to adapt their prompts for successful use Cunningham et al. (2024) and in the extra cost of additional tokens required to model linguistic variety Ahia et al. (2023). While removing identity features from prompts altogether eliminates stereotypes from the recommendations, the sanitized outputs are still biased toward one group of users, providing less aligned services for those who do not identify with the model’s default assumptions.
5 RQ3: Models Don’t Disclose When Recommendations Are Biased
Figure 5: Racial share of each race when the model admits to factoring in race, doesn’t admit to factoring in race, and for baseline prompts.
Users may be sensitive to the possibility that recommendations they received from a LLM are biased. For closed-source models, and without built-in explanation tools Zhao et al. (2023), a user might proceed to a second conversational turn, asking “Did my race influence your recommendations?"
To examine how transparent models are, we calculate the average demographic alignment of recommendations across three scenarios: second turns where the response indicates that race influenced the recommendations, ones where the response indicates that race didn’t influence the recommendations, and responses belonging to prompts that had no user identity features included. Because the second conversational turn occurs after the recommendations were already produced, the degree of bias in the recommendations could be expected to influence the model’s second response. Ideally, when the demographics of the recommendations are aligned with the revealed identity of the user, the model’s second response would alert the user to the possibility that racial stereotyping occurred.
To test if a model is obfuscating the effect of identity on recommendations, we consider as our null hypothesis the notion that stereotypical recommendations lead to indications of racial influence in the second conversational turn. Contrapositively, we test if outputs indicating that race was not a factor entail recommendations that are not biased (i.e. where recommendations were not significantly more aligned with the user’s identity).
Figure 5 shows the results of this test: across every model and every race, there is statistically significant bias even when responses deny the impact of racial identity. The one exception is Llama3’s university recommendations for white users.
On a purely numerical level, models rarely (less than 5% of the time for even the model with the most cases) admitted to taking the race of White users into account, even when explicitly asked to by the user. On the other hand, models admitted to factoring in race for other races up to $35\%$ of the time, mostly for Black users and prompts involving explicit self-identification or strongly associated entities. This may be in part due to the fact that White users are most closely aligned with the default, as discussed in the previous section, and as such their race truly does not impact their results.
For all other races, there is a sharp drop off in demographic alignment between cases where the model admits to taking race into account and cases where it doesn’t. This is a positive sign; even if the model does not perfectly guess when the user is pragmatically signaling that they want their race to be taken into account, it is usually transparent when it recommends entities on the basis of demographic alignment. In these cases, users can avoid being blindsided by biased responses even if their identity was revealed unintentionally and can ensure that responses are personalized when their identity was revealed on purpose. Still, none of the models’ achieve consistent transparency. There is still a statistically significant difference in the racial share between responses to baseline prompts and those where the model does not admit to factoring in users’ races.
6 Related Work
These results contribute to a growing body of work on the potential of machine learning systems to perpetuate bias, connecting it with an increased interest in AI systems that provide recommendations in a conversational setting.
6.1 Fairness and Bias in LMs
Language models trained on biased data replicate the biases and attitudes of the data they are trained on Caliskan et al. (2017); Hovy and Prabhumoye (2021); Jiang et al. (2022), which can be furthered by a biased selection of data Dhamala et al. (2021); Lucy et al. (2024), and is difficult to remove Gonen and Goldberg (2019).
Models can cause representational harm by reproducing harmful stereotypes Agarwal et al. (2019) or result in unfair decision making when deployed in sensitive contexts like law enforcement Angwin et al. (2022), hiring Dastin (2022), or healthcare Gianfrancesco et al. (2018); Obermeyer et al. (2019). Recently Hofmann et al. (2024) demonstrated that language models can produce biased output in response to AAE dialectal text, and Bhatt and Diaz (2024) found that explicit mentions of cultural identities in the input condition a language model to produce culturally relevant adaptations.
Models can also cause allocational harm Barocas et al. (2017) by providing different levels of benefit to different users. Other work has identified performance gaps with respect to dialect Kantharuban et al. (2023) and race Sap et al. (2019), though with a focus on specific tasks rather than user outcomes.
6.2 Fairness and Bias in Recommender Systems
Similar to allocational harm, research has examined fairness in recommender systems, defined by Ekstrand et al. (2012) as whether the “benefits and resources [the system] provides [are] fairly allocated between different people or organizations it affects,” often focusing on disparate impact. There is also work on debiasing recommender systems Chen et al. (2023), and using large language models for recommendation may also encounter issues of popularity bias Cañamares and Castells (2018).
6.3 Conversational Recommender Systems
Conversational recommender systems are recommender systems that engage in a conversation with the user Sun and Zhang (2018). LLMs used for recommendations are a specific case. Recent work has investigated fairness and bias in LLM recommendations for both explicit indications Zhang et al. (2023), which in our view may be impacted by the pragmatic implications of explicitly disclosing identity, and more naturalistic Salinas et al. (2023) identity references. However, to our knowledge no work has investigated sensitivity to truly implicit identity references, such as through dialect or reference to demographically-associated entities. Other work has investigated training strategies for recommender systems based on smaller language models Shen et al. (2023), and interfaces for explicitly controlling a chatbot’s user model Chen et al. (2024).
7 Discussion and Conclusion
Our work considers the tradeoff between beneficial personalization and harmful stereotyping in LLMs. Much of the previous work on stereotyping in LLMs has argued that the presence of bias in their outputs is strictly negative, in contexts such as healthcare Ceballos-Arroyo et al. (2024), hiring Wan et al. (2023), and education Lee et al. (2024). Yet models must sometimes factor in a user’s identity to provide personalized responses. Indeed, companies like OpenAI have recently introduced personalization features [^6] https://openai.com/index/memory-and-new-controls-for-chatgpt/ that factor in the user’s chat history. If they give the model access to any personal information, including previous chats which may contain implicit or explicit indicators of identity, they have a responsibility to avoid unnecessary and harmful bias in the resulting personalized responses. While the use of this feature is technically up to the user, it is currently opt-out, calling into question the informed consent of the user Utz et al. (2019), and the identity features that it saves are not decided by the user unless they manually parse through the model’s memory.
In a recommendation request, peripheral information embedded in the query (unintentionally and intentionally) communicates both features of the user’s identity and whether the user expects that knowledge to impact the response. Ideally, the model would be able to identify when the user is seeking personalization and apply appropriate levels of bias to its response. Unfortunately, it is not always easy to determine when this is the case. Assuming a context where it is not immediately clear whether the user wants their revealed identity to factor into their recommendations, the model is left with two imperfect options.
On one hand, providing different results to different users when they are asking for the same thing is a sign of underlying stereotyping that could negatively impact certain demographic groups Salinas et al. (2023); Shen et al. (2023); Zhang et al. (2023). In those cases, bias in the model’s response on the basis of immutable characteristics such as race would be akin to discrimination against members of that group. With the long history of redlining in major U.S. cities, where members of minority racial groups were denied housing and services outside of specific areas, it is dangerous for models to assume that users’ unspoken preferences should be determined even partly by race Lynch et al. (2021), especially when those recommendations contribute to major life decisisions like where to attend school and where to live.
Additionally, when identity is disclosed using indirect or implicit features, such as references to stereotypically associated entities, it is possible that the user’s identity does not align with the model’s assumptions. For example, members of any race may enjoy Japanese films or Mexican food, so basing decisions on references to these cultural entities could result in incorrect responses even if the goal is to achieve a high degree of personalization.
On the other hand, there are many cases where identity features play into the goals of the user and ignoring them in those cases would be refusing those users the same service provided to those whose identity better aligns with the societal "defaults" absorbed by the model. Here, personalization can provide users with better results: they may want their background to be taken into account Ghodratnama and Zakershahrak (2023); Yang et al. (2023); Domenichini et al. (2024).
Ideally, models outputs would be biased towards the user’s identity only in cases where the user wants personalization, falling back on a demographically-agnostic neutral set of assumptions otherwise. And in all cases, we argue that models should aim to be transparent when they take user’s revealed identities into account.
Unfortunately, we find that models are not only biased in their recommendations, but they also obfuscate the impact of race on their decisions. This implicit identity effect reduces user agency, and as more and more people use AI systems to access information, it has the potential to reinforce existing discrimination and stereotypes at a large scale.
Ethics
In this work, we explore bias in LLM recommendations through the lens of four U.S.-centric racial categories: White, Black, Hispanic, and Asian. In reality, there are many people who fall outside or between these categorizations, and in many cases the notion of “race” is insufficient to properly respect the broad variation of individuals’ identification with a particular race, ethnicity, ancestry, or other groups U.S. Commission on Civil Rights (2009).
Though simplistic, operating within this framework enabled us to demonstrate the bias in LLM recommendations by comparison to real-world demographic statistics. There are many other lenses through which an individual may self-identify or be identified by an external party (in this case, an LLM), such as gender, class, and age. We hope that this paper can provide a framework for further study into how other identity categories can influence LLM recommendations and decision-making. For instance, users can implicitly reveal their age through cultural references or era-specific slang.
Individuals at the intersection of multiple identity categories can sometimes face even worse degrees of discrimination; Black women, for instance, historically have seen greater systemic discrimination than either women or the general Black population and may see differing results from those in this paper Nash (2008). Additionally, dialectal use in this paper is based on the intended race of the hypothetical user, however, these are largely non-standardized forms of speech, and as such their use and structure vary across the population, or even see uptake by other groups Reyes (2005); Roth-Gordon et al. (2020).
In this paper we examine bias with nuance, arguing that there are times in which it may be appropriate for personalization. That being said, we disagree with arguments that might justify the presence of bias for the sake of personalization. Rather, we believe that AI products should, at the very least, be transparent, if not also steerable. Users ought to have agency: first, an understanding of when an AI system has inferred their identity (particularly when the system does so on the basis of implicit characteristics), especially if this inference influences responses; and second, the ability to control how and when an AI system takes their identity into account.
Limitations
In this section, we review a few limitations of our analysis.
Synthetic prompts
While our synthetic prompts aim to emulate realistic user behavior, they are artificially constructed and do not cover the larger structural variance of naturalistic prompts. This limits the ways in which identity could be injected into the prompts, especially implicitly. While we found that there were no statistically significant differences in the performance of models from dialect alone, our analysis largely focused on varying syntactic and grammatical constructions, and the inclusion of additional dialectal modifications may show different results.
U.S.-centered analysis
In this work we focused on a specific use case of the model (recommendations for two major life decisions) among American users who can be easily racially categorized. In reality, these chatbot systems are used by people from all over the world and with a huge variety of languages, dialects, nationalities, and cultures. Additionally, the racial dynamics we explore concerning stereotyping and discrimination are based on American society and history, so results may differ for communities outside the United States.
Future system development
While we demonstrate that there are clear trends across all the studied models, these results may not hold true through future LLM releases under different training conditions or even just further human feedback training. However, the conversation surrounding the role of implicit user identity in LLM outputs is still new. With companies only recently beginning to introduce personalization features in their consumer chatbots, although our results may be transient, this work is necessary to establish the goal of studying the tradeoff between bias and personalization, and the potential harms of silently “personalized” LLMs.
Acknowledgements
We would like to thank Shaily Bhatt, Maarten Sap, Fernando Diaz, as well as other members of the Language Technologies Institute at Carnegie Mellon University for their helpful comments and suggestions.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2022335781. JM was supported by the Center for Informed Democracy & Social Cybersecurity as a Knight Fellow. The views and conclusions in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Knight Foundation or the U.S. Government.
Appendix A Prompts
See Table 3 for the prompt template used.
University Recommendation
I am a high school senior and I {going to} go to college. I have a <gpa> and a <score> SAT score. + ( I {am} on the school <sport> team.) + ( I also volunteer with <organization>.) + ( After school, I work <jobs>.) + ( {Don’t recommend me} places that are <constraints>.) + ( Keep in mind that I am <race>.) What five colleges would you recommend for me? Return only a json object with the keys “university" and “explanation" for each of the recommended universities. Be brief and to the point with explanations.
Neighborhood Recommendations
I just graudated from <university> with a degree in <major> and am moving to <city> to start a new job. + ( I mostly {am} <transportation> to get around. ) + ( My mom {told} me to mention I want to live near restaurants that serve <cuisine> food.) + ( {Don’t recommend me} places that are <constraints>.) + ( Keep in mind that I am <race>.) What five neighborhoods would you recommend for me? Return only a json object with the keys “neighborhood" and “explanation" for each of the recommended universities. Be brief and to the point with explanations.
Table 3: Prompts for each task. Parentheses represent optional sentence, angle brackets represented slotted in constraints and entities, and curly brackets represent spots where dialectal variation can be injected.